Winston Industries had sales of $843,800 and costs of $609,900. The firm paid $38,200 in interest and $18,000 in dividends. It also increased retained earnings by $62,138 for the year. The depreciation was $76,400. What is the average tax rate? O 32.83 percent O 33.33 percent 38.17 percent O 43.39 percent
Answer #1

net income for the year

=increase in retained earnings+dividends

=62138+18000

=80138

net income before taxes

=sales-costs-interest costs-depreciation

=843800-609900-38200-76400

=119300

taxes=net income before taxes-net income for the year

=119300-80138

=39162

average tax rate is=taxes/net income before taxes

=39162/119300

=32.83%

the above is answer..